What is the length of the diagonal of a square with area 40.5 square inches
Found 2 solutions by nyc_function, stanbon:
Answer by nyc_function(2741)

Here's what to do:
1-E-mail your question to me at nycfunction@yahoo.com
2-Include your fax number (most people have a fax number at home these days)
3-I will then fax the entire question and fully explained answer.
NYC Math Tutor

Answer by stanbon(75887)

What is the length of the diagonal of a square with area 40.5 square inches
---
side^2 = 40.5
side = sqrt(40.5)
----
Use Pythagoras
---
diagonal = sqrt(side^2 + side^2)
d = sqrt(40.5 + 40.5)
d = sqrt(2(40.5))
d = sqrt(81)
d = 9 inches
====================
cheers,
Stan H.